# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ III.1 Description of $\mathbb{Z}_{4}$
For $N=4$ the explicit matrix representations at a given site are:
@@ III.1.1 Bulk
where H.C. denotes the hermitian conjugate and the sum is over vertices and plaquettes belonging to the bulk region. Since all the terms of the Hamiltonian commute, they define a stabilizer code generated by all the star and plaquette operators:
@@
When two anyons come close together, they can be regarded as one emergent anyon; this process is called fusion. In Abelian theories, fusion outcomes are deterministic, which means the fusion of any two anyons results in only one new anyon. Explicitly, we have:
@@
As is common in topological phases, bosonic excitations (anyons with topological spin =1=1) can still braid non-trivially with other excitations. The braiding of two anyons $a$ and $b$ around each other produces a $U(1)$ phase denoted here by $B(a,b)$. In Abelian theories, this phase can be directly computed from the $\mathcal{T}$-matrix as:
@@
The braiding statistics of anyons play a crucial role in defining the $\mathbb{Z}_{4}$ models as quantum error-correcting surface, toric, or other related codes. Closed Wilson loops of anyons commute with all the Hamiltonian terms and thus define/implement logical operators acting within the code space. The caveat is that if two loops of anyons braid non-trivially, then their actions are dependent. In addition, loops that are deformed into one another, via the action of the terms in the Hamiltonian, are also equivalent.
III.1.2 External Boundaries
@@
In this section, we focus on the generalization of the $\mathbb{Z}_{2}$ case to the gapped boundaries of TQFTs described by (twisted) quantum doubles based on the $D(\mathbb{Z}_{4})$. Note that our notion of boundaries, thus far, corresponds to a boundary between a non-trivial TQFT and the topologically trivial vacuum, where no anyons exist. This means that at the bulk-to-boundary map, a bulk deconfined (an anyon that does not need extra energy to propagate) anyon may only condense (be identified with the vacuum) or become confined depending on the details of the boundary. The confined anyons will still be gapped, and they will live on the (1+1)(1+1)D boundary. In fact, the (1+1)(1+1)D gapped boundaries of topological phases are completely determined by the anyons they condense [63, 46]. These anyons for a certain boundary form a subgroup of the fusion group, which is called a Lagrangian subgroup. This subgroup $\mathcal{L}$ satisfies three properties:
@@
Intuitively, the Lagrangian subgroup describes a group of anyons that can be consistently identified with the vacuum. In addition, identifying this group with a vacuum will confine all other anyons outside the subgroup. From the spin and braiding statistics of the $\mathbb{Z}_{4}$ model, we can find three boundaries:
@@
Suppose the model lives on a square surface $\Sigma$ with a boundary $\partial\Sigma$ Fig. 2. Then the Hamiltonian in Eq. (6) has to be changed to include boundary terms in order to realize these three boundaries:
$H_{\mathbb{Z}_{4},\text{Smooth}}$ $=H_{\mathbb{Z}_{4}}-\sum_{v\in\partial\Sigma}(A(v))+H.C.$~ (14)
$H_{\mathbb{Z}_{4},\text{Rough}}$ $=H_{\mathbb{Z}_{4}}-\sum_{p\in\partial\Sigma}(B(p))+H.C.$~
$H_{\mathbb{Z}_{4},\text{Even}}$ $=H_{\mathbb{Z}_{4}}-\sum_{p,e\in\partial\Sigma}(B^{2}(p)+X^{2}(e))+H.C.$
@@
We see that the boundary terms are derived from the bulk terms with possibly a higher power or restricted support. We will formalize this notion in Sec.IV, where a general algorithm is given that constructs all boundaries of Abelian twisted quantum doubles in terms of Pauli stabilizers. A $\mathbb{Z}_{4}$ surface code will then be a choice of these boundaries. We provide one example here, while other examples can be treated analogously using methods of V.
@@
The ground state is two-fold degenerate. We discuss methods for computing the ground state degeneracy ($\operatorname{GSD}$) for different surfaces and boundaries in Sec. V. The logical operators can be chosen to be the vertical world-line of a $e^{2}$ that condenses on the even and the rough boundaries, in addition to the horizontal $m$ world-line string that condenses on the smooth boundaries. From the braiding equation Eq. (12), we can read $B(e^{2},m)=-1$, and from the fusion Eq. (9) we have the relations $(e^{2})^{2}=1$. The square of the entire $m$ string condenses on the top even boundary, and it is therefore identified with the identity. Since the two logical operators anticommute and square to identity, they generate the one-qubit Pauli group, and we define the $e^{2}$ and $m$ worldline strings as the $\bar{Z}$ and $\bar{X}$ logical operators, respectively. Note that, had the top boundary been rough, rather than even, the code would have defined a logical $d=4$ qudit.
III.1.3 Domain walls and defects
In addition to boundaries with vacuum, topological phases can have boundaries with other compatible phases or, in particular, with the same phase. Those boundaries are called domain walls for distinction. They are also called defects, as they usually correspond to defects in the lattice. We reserve the term defects for the 0D ends of a finite domain wall. Here, we focus on internal invertible domain walls within $D(\mathbb{Z}_{4})$. An invertible domain wall is a wall that can be followed by another wall, producing the transparent (trivial) domain wall that acts as identity. In $D(\mathbb{Z}_{4})$, we have four invertible walls. Labeled by their action on the generators $e$ and $m$, they are:
@@ III.1.3 Domain walls and defects
The exact layout of the stabilizers implementing the four domain walls is shown in Fig. 3. At the end of a finite domain wall, a $0D$ defect lives. These defects can be used for quantum computing on their own [53, 65, 66, 67, 68, 69, 47]. Sec. IV discusses how domain walls, in addition to $0D$ defects, can be constructed systematically from bulk stabilizers. Note that the domain walls in Fig. 3 realize their mappings in any $\mathbb{Z}_{n}$ for appropriate generalized Pauli matrices, Eq.(3) (but they do not necessarily exhaust all the invertible domain walls for $\mathbb{Z}_{n}$ with $n>4$).
@@
Figure 3: Stabilizers layout for the 33 non-trivial invertible $\mathbb{Z}_{4}$ domain walls ending with 0D defects. Definitions of symbols are given in Eqs. (17) and (18).
III.1.4 Topological Codes
Figure 4: Left: a two-logical-qudit $\mathbb{Z}_{4}$ surface code with $N=3$ smooth punctures. The logical $\widetilde{Z}$ operators can be taken to be $e$ loops around N1N-1 holes. The logical $\widetilde{X}$ operators are $m$ strings from the N1N-1 holes to the first one. Right: another two-logical-qudit $\mathbb{Z}_{4}$ surface code with $N=3$ ($e\leftrightarrow m$) twists. The logical $\widetilde{Z}$ operators can be taken to be $e$ loops around N1N-1 twists. The logical $\widetilde{X}$ operators are hybrid $m$-$e$ loops between the N1N-1 twists and the first one.
After discussing the bulk, boundaries, and domain walls of $D(\mathbb{Z}_{4})$, we present illustrative examples on how these components can be put together to construct different topological codes in this section. These constructions include, but are not limited to, adding punctures to increase the number of qubits in $\mathbb{Z}_{2}$ surface code [3, 70, 71], adding twists that affect the $\operatorname{GSD}$ [53, 72, 73].
Punctures–Left panel in Fig. 4 illustrates the case with $3$ smooth punctures. Choosing different boundary conditions can increase $\operatorname{GSD}$, here, we assume that the model lives on a sphere or has an outer rough boundary for simplicity. The logical operators are the two $m$ strings between the smooth punctures. We take these to represent the $\widetilde{X}_{1}$ and $\widetilde{X}_{2}$ logical Pauli operators. While the two $e$ loops around the holes represent generalized Paulis: $\widetilde{Z}_{1}$ and $\widetilde{Z}_{2}$. This is evident from the braiding of $B(e,m)=i$ Eq. (12). Further, since the bulk is $\mathbb{Z}_{4}$, the fourth power of each of them equals identity, and we have two $4$-qudits. In general, we have for $N$-holes of this type:
$\operatorname{GSD}_{\text{N-smooth/rough holes}}=4^{N-1}$ (19)
We discuss in more detail how the ground state degeneracy can be calculated exactly in V.
Twists–Similarly, we can introduce defects that correspond to the domain wall $W^{2}:(e,m)\mapsto(m,e)$ of Fig. 4. For simplicity, the model can be taken to have rough boundary conditions (or smooth but not both), or equivalently, live on a sphere. One can form a loop of $e$ (alternatively, $m$ anyons) around the whole domain wall. In the presence of more than one twist, such loops will not be contractible and can be taken to represent Pauli $\widetilde{Z}_{1}$, $\widetilde{Z}_{2}$ operators. As before, their fourth power is identity by virtue of the $\mathbb{Z}_{4}$ bulk. In addition, an $e$ and an $m$ particle coming from the two opposite sides of one twist can condense. From another perspective, it can be viewed as the permutation of the $e$ into an $m$ particle. This introduces new loops that are half $e$ and half $m$, which are drawn in blue in Fig. 4. The braiding $B(e,m)=i$ again shows that these loops represent the $\widetilde{X}_{1}$, $\widetilde{X}_{2}$ operators. This generalizes the twists given in Ref. 53. For $N$ twists of this type, we have:
$\operatorname{GSD}_{\text{N-twists}}=4^{N-1}$ (20)
Figure 5: A four-logical-qubit $\mathbb{Z}_{4}$ surface code with $N=3$ even punctures. The logical $\widetilde{Z}$ operators can be taken to be $e$ and $m$ loops around N1N-1 holes. The logical $\widetilde{X}$ operators are $m^{2}$ or $e^{2}$ strings from the N1N-1 holes to the first one. In general, we have 2(N1)2(N-1) qubits.
New Punctures–In addition to these codes that generalize $D(\mathbb{Z}_{2})$ constructions, new phenomena can arise when considering the new even boundary of the $D(\mathbb{Z}_{4})$. Let us take the $D(Z_{4})$ surface code with smooth external boundaries (or rough or on a sphere). Next, we poke three holes with the even boundary that condenses $\mathcal{L}_{\text{Even}}=\{1,e^{2},m^{2},e^{2}m^{2}\}$. In this case, we can transport any anyon in this Lagrangian subgroup between two boundaries. With only one hole, we have $\operatorname{GSD}=1$ as expected. However, with two holes we gain $2$ qubits, not a qubit like the $D(\mathbb{Z}_{2})$ and not a 4-dit like the smooth or rough holes case. In general, we have $2N$ qubits for $N$ holes with even boundaries. More concretely, the logical $\widetilde{X_{1}}$ can be chosen to be a string of $X^{2}$ operators transporting an $m^{2}$ anyon between the two holes. The $\widetilde{X_{2}}$ logical operator will be a string of $Z^{2}$ operators transporting an $e^{2}$ anyon. These two anyons commute Eq. 12 thus we have $\widetilde{X_{1}},\widetilde{X_{2}}]=0$. The logical $\widetilde{Z_{1}}$ operator will be a loop of Pauli $Z$ around one hole, while the logical $\widetilde{Z_{2}}$ will be an $X$ loop around the same hole. The two loops commute with each other as they cross an even number of times. Further, each loop anti-commutes with its logical $\widetilde{X}$ operator. We see that interestingly, the logical $\widetilde{X}$ error rate is doubled in comparison to the logical $\widetilde{Z}$ error rate. Even though the degeneracy is the same as other holes.
$\operatorname{GSD}_{\text{N-even holes}}=4^{N-1}$ (21)
III.2 Double Semion Model
@@
We are interested in the error-correcting properties of the $\mathbb{Z}_{4}$ phase and its closely related phases. In this section, we review the doubled semion (DS) phase, which is closely related to $\mathbb{Z}_{4}$ as we will discuss later in Sec. IV. The doubled semion (DS) phase is an Abelian topological order first realized as a string-net model [74]. It is a twisted $\mathbb{Z}_{2}$ gauge theory. It is also the simplest twisted quantum double, which is a generalization of the Kitaev quantum double [17]. All twisted quantum doubles with Abelian orders can be described in terms of Pauli stabilizers [40].
III.2.1 Bulk
@@
To begin, let us consider the bulk stabilizer Hamiltonian describing the DS phase as described in Ref. 40. As in the $D(\mathbb{Z}_{4})$ case, the model will live on a lattice consisting of qudits. As illustrated in Eq. (22), the operator $F(v,p)=A(v)B(p)$ denotes the product of adjacent $\mathbb{Z}_{4}$ vertex and plaquette operators, while the $C$ operators are a generalization of the $\mathbb{Z}_{2}$ twist [72]. However, in the context of $\mathbb{Z}_{4}$, since $X^{2},Z^{2}]=0$ these operators are now mutually compatible and they may overlap.
@@
This redundancy in the ribbon operators stems from the stabilizer embedding of DS in $\mathbb{Z}_{4}$ as discussed in IV. In this embedding, the ribbon operator for the $e^{2}m^{2}$ anyons in $\mathbb{Z}_{4}$ commutes with all the stabilizers of the Hamiltonian. One choice for the ribbon operators for the different anyons is shown in Fig. 6. Multiplying the ribbon operators by the one for $e^{2}m^{2}$ results in another equivalent ribbon. In this code, the $F$ operators act as syndromes detecting the boson while the $B$ and $C$ operators, having eigenvalue 1-1 denote the presence of a (anti-)semion. Reference 40 showed how this phase may be put on a torus to construct a pair of logical qubits that have logical operators of weight two. In the next sections, we will generalize this result.
III.2.2 External Boundaries
As we briefly described in Sec. III.1.2 and will expand on in Sec. IV, each boundary is completely determined by its Lagrangian subgroup III.1.2. In the case of DS, we only have one non-trivial boson $b$ and thus only one Lagrangian subgroup:
@@ III.2.2 External Boundaries
$H_{\text{DS},\text{Boundary}}$ $=H_{\text{DS}}-\sum_{v,e\in\partial\Sigma}(A^{2}(v)+Z^{2}(e))$ (30)

### v4

@@ III.1 The $D(\mathbb{Z}_{4})$ Quantum Double
For $N=4$, the explicit matrix representations at a given site are:
@@ III.1.1 $D(\mathbb{Z}_{4})$ Bulk
where H.C. denotes the hermitian conjugate and the sum is over vertices and plaquettes belonging to the bulk region.
[Uncaptioned image] (6)
Since all the terms of the Hamiltonian commute, they define a stabilizer code generated by all the star and plaquette operators:
@@
When two anyons come close together, they can be regarded as one emergent anyon; this process is called fusion [113]. In Abelian theories, fusion outcomes are deterministic, which means the fusion of any two anyons results in only one new anyon [107, 80, 178]. Explicitly, we have:
@@
As is common in topological phases, bosonic excitations (anyons with topological spin =1=1) can still braid non-trivially with other excitations. The braiding of two anyons $a$ and $b$ around each other produces a $U(1)$ phase denoted here by $B(a,b)$. In Abelian theories, this phase can be directly computed from the $\mathcal{T}$-matrix as [107]:
@@
The braiding statistics of anyons play a crucial role in defining the $\mathbb{Z}_{4}$ models as quantum error-correcting surface, toric, or other related codes. Closed Wilson loops of anyons commute with all the Hamiltonian terms and thus define/implement logical operators acting within the code space [24]. The caveat is that if two loops of anyons braid non-trivially, then their actions are dependent. In addition, loops that are deformed into one another, via the action of the terms in the Hamiltonian, are also equivalent.
III.1.2 $D(\mathbb{Z}_{4})$ Boundaries
@@
In this section, we focus on the generalization of the $\mathbb{Z}_{2}$ case to the gapped boundaries of TQFTs described by (twisted) quantum doubles based on the $D(\mathbb{Z}_{4})$. Note that our notion of boundaries, thus far, corresponds to a boundary between a non-trivial TQFT and the topologically trivial vacuum, where no anyons exist. This means that at the bulk-to-boundary map, a bulk deconfined (an anyon that does not need extra energy to propagate) anyon may only condense (be identified with the vacuum) or become confined depending on the details of the boundary. The confined anyons will still be gapped, and they will live on the (1+1)(1+1)D boundary. In fact, the (1+1)(1+1)D gapped boundaries of topological phases are completely determined by the anyons they condense [103, 114]. These anyons for a certain boundary form a subgroup of the fusion group, which is called a Lagrangian subgroup. This subgroup $\mathcal{L}$ satisfies three properties [67, 178, 103]:
@@
Intuitively, the Lagrangian subgroup describes a group of anyons that can be consistently identified with the vacuum. In addition, identifying this group with a vacuum will confine all other anyons outside the subgroup [126]. From the spin and braiding statistics of the $\mathbb{Z}_{4}$ model, we can find three boundaries:
@@
Suppose the model lives on a square surface $\mathcal{M}$ with a boundary, $\partial\mathcal{M}$ Fig. 2. Then the Hamiltonian in Eq. (5) has to be changed to include boundary terms in order to realize these three boundaries:
$H_{\mathbb{Z}_{4},\text{Smooth}}$ $=H_{\mathbb{Z}_{4}}-\sum_{v\in\partial\mathcal{M}}(A(v))+H.C.\penalty 10000\$(14)
$H_{\mathbb{Z}_{4},\text{Rough}}$ $=H_{\mathbb{Z}_{4}}-\sum_{p\in\partial\mathcal{M}}(B(p))+H.C.\penalty 10000\$
$H_{\mathbb{Z}_{4},\text{Even}}$ $=H_{\mathbb{Z}_{4}}-\sum_{p,e\in\partial\mathcal{M}}(B^{2}(p)+X^{2}(e))+H.C.$
@@
We see that the boundary terms are derived from the bulk terms with possibly a higher power or restricted support. We will formalize this notion in Sec.V, where a general algorithm is given that constructs all boundaries of Abelian twisted quantum doubles in terms of Pauli stabilizers. A $\mathbb{Z}_{4}$ surface code will then be a choice of these boundaries. We provide one example here, while other examples can be treated analogously using methods of VI.
@@
The ground state is two-fold degenerate. We discuss methods for computing the ground state degeneracy ($\operatorname{GSD}$) for different surfaces and boundaries in Sec. VI. The logical operators can be chosen to be the vertical world-line of a $e^{2}$ that condenses on the even and the rough boundaries, in addition to the horizontal $m$ world-line string that condenses on the smooth boundaries. From the braiding equation Eq. (12), we can read $B(e^{2},m)=-1$, and from the fusion Eq. (9) we have the relations $(e^{2})^{2}=1$. The square of the entire $m$ string condenses on the top even boundary, and it is therefore identified with the identity. Since the two logical operators anticommute and square to identity, they generate the one-qubit Pauli group, and we define the $e^{2}$ and $m$ worldline strings as the $\bar{Z}$ and $\bar{X}$ logical operators, respectively. Note that, had the top boundary been rough, rather than even, the code would have defined a logical $d=4$ qudit.
III.1.3 $D(\mathbb{Z}_{4})$ Domain walls and defects
In addition to boundaries with vacuum, topological phases can have boundaries with other compatible phases or, in particular, with the same phase. Those boundaries are called domain walls for distinction. They are also called defects, as they usually correspond to defects in the lattice [21]. We reserve the term defects for the 0D ends of a finite domain wall [11, 10, 9]. Here, we focus on invertible domain walls within $D(\mathbb{Z}_{4})$. An invertible domain wall is a wall that can be followed by another wall, producing the transparent (trivial) domain wall that acts as identity [113]. In $D(\mathbb{Z}_{4})$, we have four invertible walls. Labeled by their action on the generators $e$ and $m$, they are:
@@ III.1.3 $D(\mathbb{Z}_{4})$ Domain walls and defects
The exact layout of the stabilizers implementing the four domain walls is shown in Fig. 3. At the end of a finite domain wall, a $0D$ defect lives. These defects can be used for quantum computing on their own [21, 22, 187, 104, 164, 149, 52, 53, 111, 110]. Sec. V discusses how domain walls, in addition to $0D$ defects, can be constructed systematically from bulk stabilizers. Note that the domain walls in Fig. 3 realize their mappings in any $\mathbb{Z}_{n}$ for appropriate generalized Pauli matrices, Eq.(3) (but they do not necessarily exhaust all the invertible domain walls for $\mathbb{Z}_{n}$ with $n>4$).
@@
Figure 3: Stabilizers layout for the $3$ non-trivial invertible $\mathbb{Z}_{4}$ domain walls ending with 0D defects. Definitions of symbols are given in Eqs. (17) and (18).
III.2 Double Semion Twisted Quantum Double
@@
We are interested in the error-correcting properties of the $\mathbb{Z}_{4}$ phase and its closely related phases. In this section, we review the doubled semion (DS) phase, which is closely related to $\mathbb{Z}_{4}$ as we will discuss later in Sec. V. The doubled semion (DS) phase is an Abelian topological order first realized as a string-net model [125]. It is a twisted $\mathbb{Z}_{2}$ gauge theory. It is also the simplest twisted quantum double, which is a generalization of the Kitaev quantum double [92]. All twisted quantum doubles with Abelian orders can be described in terms of Pauli stabilizers [66].
III.2.1 DS Bulk
@@
To begin, let us consider the bulk stabilizer Hamiltonian describing the DS phase as described in Ref. 66. As in the $D(\mathbb{Z}_{4})$ case, the model will live on a lattice consisting of qudits. As illustrated in Eq. (19), the operator $F(v,p)=A(v)B(p)$ denotes the product of adjacent $\mathbb{Z}_{4}$ vertex and plaquette operators, while the $C$ operators are a generalization of the $\mathbb{Z}_{2}$ twist [117]. However, in the context of $\mathbb{Z}_{4}$, since $X^{2},Z^{2}]=0$ these operators are now mutually compatible and they may overlap.
@@
This redundancy in the ribbon operators stems from the stabilizer embedding of DS in $\mathbb{Z}_{4}$ as discussed in V. In this embedding, the ribbon operator for the $e^{2}m^{2}$ anyons in $\mathbb{Z}_{4}$ commutes with all the stabilizers of the Hamiltonian. One choice for the ribbon operators for the different anyons is shown in Fig. 4. Multiplying the ribbon operators by the one for $e^{2}m^{2}$ results in another equivalent ribbon. In this code, the $F$ operators act as syndromes detecting the boson while the $B$ and $C$ operators, having eigenvalue 1-1 denote the presence of a (anti-)semion. Reference 66 showed how this phase may be put on a torus to construct a pair of logical qubits. In the next sections, we will generalize this result.
III.2.2 DS Boundaries
As we briefly described in Sec. III.1.2 and will expand on in Sec. V, each boundary is completely determined by its Lagrangian subgroup III.1.2. In the case of DS, we only have one non-trivial boson $b$ and thus only one Lagrangian subgroup:
@@ III.2.2 DS Boundaries
$H_{\text{DS},\text{Boundary}}$ $=H_{\text{DS}}-\sum_{v,e\in\partial\mathcal{M}}(A^{2}(v)+Z^{2}(e))$ (27)